OpenStudy (anonymous):

Solve this system \(\large (\sqrt [ 3 ]{ 15x^{ 6 }-9y^{ 4 }+7y^{ 2 }-16x+4 } )^{ 2013 }\\\large+(\sqrt [ 4 ]{ 20y^{ 6 }-8x^{ 2 }+7y^{ 6 }-17y^{ 2 }-2 } )^{ 2014 }\\\large =(\sqrt [ 5 ]{ 5x^{ 6 }+6x^{ 5 }-4y^{ 6 }-3y^{ 3 }-3y-4x+3 } )^{ 2015 }\) and \(\large 9(x^{10}-10x^9+50x^8-160x^7+360x^6+719x^4-636x^3+392x^2\\\large-152x+31) (81y^5-270y^4+360y^3-243y^2+84y-11)^2\\\large =x^ 6-6x^ 5+18x^ 4y+6x^ 4-72x^3y+16x^3+108x^2y^2-12x^2\\\large-216xy^2 +144xy-24x+216y^3-21y^2+72y-8\)

OpenStudy (vishweshshrimali5):

My god !! @JungHyunRan it would take me a minute at least to read it. :D

OpenStudy (anonymous):

^-^

OpenStudy (vishweshshrimali5):

Well give me some time and I would most probably come up with a solution. I am ill today so cannot be too sure of myself ;)

OpenStudy (anonymous):

Just wait...

OpenStudy (anonymous):

I am just calling Einstein, may be He can do it... :P

OpenStudy (anonymous):

lol @waterineyes Einstein???

OpenStudy (vishweshshrimali5):

Yeah a close friend of @waterineyes ;)

OpenStudy (vishweshshrimali5):

Okay one way is to use some online equation solver :D

OpenStudy (anonymous):

When I asked that @JungHyunRan has posted a question on OS, he just cut the call... :P

OpenStudy (anonymous):

haha... so funny :P

OpenStudy (anonymous):

May be he knows how difficult a question you can ask... :)

OpenStudy (anonymous):

What happened @vishweshshrimali5 I think you were typing something... :P

OpenStudy (vishweshshrimali5):

Yeah internet problem. I am planning to have a break up with internet :'( It does not let me check my mail :P

OpenStudy (anonymous):

I am missing @mukushla here... :(

OpenStudy (vishweshshrimali5):

Same here

OpenStudy (vishweshshrimali5):

Well let me concentrate XD :P

ganeshie8 (ganeshie8):

this problem doesn't look like @mukushla 's level, it just looks like a pain to process the long equations and bit boring :o

OpenStudy (vishweshshrimali5):

May be collecting like terms may help for second equation

OpenStudy (vishweshshrimali5):

I am really cursing the person who created this problem. Couldn't think of anything short and simple to read and solve ?

OpenStudy (vishweshshrimali5):

\[\large (\sqrt [ 3 ]{ 15x^{ 6 }-9y^{ 4 }+7y^{ 2 }-16x+4 } )^{ 2013 }\\\large+(\sqrt [ 4 ]{ 20y^{ 6 }-8x^{ 2 }+7y^{ 6 }-17y^{ 2 }-2 } )^{ 2014 }\\\large =(\sqrt [ 5 ]{ 5x^{ 6 }+6x^{ 5 }-4y^{ 6 }-3y^{ 3 }-3y-4x+3 } )^{ 2015 }\] 2013 is divisible by 3 !!! \[\large ({ 15x^{ 6 }-9y^{ 4 }+7y^{ 2 }-16x+4 } )^{ 671 }\\\large+(\sqrt [ 2 ]{ 20y^{ 6 }-8x^{ 2 }+7y^{ 6 }-17y^{ 2 }-2 } )^{ 1007 }\\\large =({ 5x^{ 6 }+6x^{ 5 }-4y^{ 6 }-3y^{ 3 }-3y-4x+3 } )^{ 403 }\]

OpenStudy (vishweshshrimali5):

Now my main purpose would be to think of something for this: \[\large{\sqrt{20y^6-8x^2+7y^6 - 17y^2 - 2}}\]

OpenStudy (vishweshshrimali5):

@JungHyunRan 20y^6 + 7y^6 are you sure ?

OpenStudy (astrophysics):

Simplify it a bit other than that I don't think you can do anything lol

OpenStudy (vishweshshrimali5):

I mean such a question creator would not be that lenient on solvers

OpenStudy (vishweshshrimali5):

Can you please recheck the equations ?

OpenStudy (vishweshshrimali5):

I am 100% sure it was something else. I have a lot of experience with such mad mathematicians :'( They don't think in favour of us solvers... Believe me

OpenStudy (vishweshshrimali5):

@ganeshie8 left giving me medal :D

Parth (parthkohli):

Looks like there is some work spotting perfect square terms.

OpenStudy (anonymous):

@vishweshshrimali5 It may be 20y^6 + 7x^6

OpenStudy (vishweshshrimali5):

No no @JungHyunRan I am pretty sure it is something else. I checked using that too but that would not give us the answer.

OpenStudy (vishweshshrimali5):

I can say this because I have solved this problem somewhere but where ??????? God know :(

OpenStudy (vishweshshrimali5):

At that time I had found the answer...

OpenStudy (vishweshshrimali5):

Please please please recheck the question ........

OpenStudy (anonymous):

This is question, but it was wrote 20y^6+7y^6 :3

OpenStudy (vishweshshrimali5):

Well if that is the case let me retry :)

OpenStudy (vishweshshrimali5):

Well first let me simplify the second equation :) My previous two tries have gone in vain

OpenStudy (vishweshshrimali5):

Well I found something similar yet different: http://diendantoanhoc.net/forum/index.php?/topic/119665-left-sqrt315x6-9y47y2-16x4-right-2013left-sqrt420y6-8x27x2-17y-2-right-2014/ I don't know what is written there but the equation looks different yet little bit problematic as it contains -8x^2+7x^2 which I think is NOT what the person really wanted to ask in the question.

OpenStudy (vishweshshrimali5):

I was not able to get any useful result as the equations are very long and complex. I would leave a comment whenever I find out any result... @JungHyunRan Till then can you please have this problem checked from the source as there are two different versions of the same question. Vishwesh

OpenStudy (anonymous):

\[(15x^6 - 16x -9y^4 +7y^2 +4)^(671) + (-8x^2 +27y^6 - 17y^2 -2)^(1007/2)= (5x^6 +6x^5 -4x -4y^6 -3y^3 - 3y +3)^(403)\] (671),(1007/2) and (403) are to the power ^

OpenStudy (anonymous):

(5x^6 + 6x^5 - 4x - 4y^6 - 3y^3 -3y +3)^(403)

OpenStudy (vishweshshrimali5):

@Mateaus , I too solved up till this step but I don't think it is reasonable to use binomial formula for expanding the expressions :)

OpenStudy (anonymous):

Yeah I see, I tried the second one and ended with over 10 expansions

OpenStudy (vishweshshrimali5):

\(\color{blue}{\text{Originally Posted by}}\) @vishweshshrimali5 \[\large (\sqrt [ 3 ]{ 15x^{ 6 }-9y^{ 4 }+7y^{ 2 }-16x+4 } )^{ 2013 }\\\large+(\sqrt [ 4 ]{ 20y^{ 6 }-8x^{ 2 }+7y^{ 6 }-17y^{ 2 }-2 } )^{ 2014 }\\\large =(\sqrt [ 5 ]{ 5x^{ 6 }+6x^{ 5 }-4y^{ 6 }-3y^{ 3 }-3y-4x+3 } )^{ 2015 }\] 2013 is divisible by 3 !!! \[\large ({ 15x^{ 6 }-9y^{ 4 }+7y^{ 2 }-16x+4 } )^{ 671 }\\\large+(\sqrt [ 2 ]{ 20y^{ 6 }-8x^{ 2 }+7y^{ 6 }-17y^{ 2 }-2 } )^{ 1007 }\\\large =({ 5x^{ 6 }+6x^{ 5 }-4y^{ 6 }-3y^{ 3 }-3y-4x+3 } )^{ 403 }\] \(\color{blue}{\text{End of Quote}}\) Though I am thinking of finding out the square root of the second term (one with the power 1007 and a square root

OpenStudy (vishweshshrimali5):

Yeah @Mateaus and Wolfram would not allow me to expand and solve or simplify them.

OpenStudy (vishweshshrimali5):

But I am surprised that a person who created this good problem would be so lenient for us to write 20y^6 + 7y^6 instead of 27y^6... It just doesn't seem correct. What do you think @Mateaus

OpenStudy (anonymous):

hmm what if 7y^6 wasn't mean to be raised to the power of 6 but instead to another.

OpenStudy (anonymous):

@vishweshshrimali5 unless the person who created it did it for a reason.

OpenStudy (vishweshshrimali5):

Yup you are correct :)